OpenStudy (anonymous):

4. A Honda with a dealer invoice price of $19,700 was retail priced at $23,000. (a) How much is the percent markup based on selling price? Round to the nearest hundredth percent. (b) How much is the markup percent if cost had been the case (use the equivalent markup formula)? Round to the nearest hundredth percent.

OpenStudy (jdoe0001):

he got the car for $19,700, and he's selling it for $23,000 what's the difference? how much more is he selling it for? well 23,000-19,700 = 3,300 how much is 3,300 in percentage from 19,700? well \(\bf \textit{n\% of m is }=\cfrac{n}{100}\cdot m\qquad thus\\ \quad \\ 3,300=\cfrac{n}{100}\cdot 19,700\implies \cfrac{3,300\cdot 100}{19,700}=n\)

OpenStudy (anonymous):

16.75% ?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

I do not understand the difference they are asking between question A and question B it sounds like the same question to me. but I suck at math.

OpenStudy (jdoe0001):

well, I don't either.... but yes is basically the same I'd say so for A) sounds they're asking for the selling price relative to the selling price... that is, if 19,700 is 100%, how much is 23,000? well, since the markup went up by 16.75%, then 23,000 is 116.75% from the original cost B) what if the cost had been the case, or I'd think, how much will it be if in relation to the original cost of 19,700? well. the markup of 3,300 is 16.75% of the original cost

OpenStudy (jdoe0001):

so they're asking for the same value, but just from different angles of view

OpenStudy (anonymous):

Ok thank you very much for your help/ ;-)

OpenStudy (jdoe0001):

yw